a charter bus travels 210 miles in 3 1/2 hours. assuming that the distance is directly proportional to the time traveled, how far will the bus travel in 6
hours

Answer by jndarrell(58)

Your proportion should set up like this: miles/hours or hours/miles. Either or, just make sure that both sides of the equation have the units in the same place (mi/hr=mi/hr not mi/hr=hr/mi)
210/3.5=x/6
Cross multiply
3.5x=1260
Divide both sides by 3.5 to get x alone
x=360
Make sure you include the units in your solution (360miles).
To check the work, plug the solution back into the original equation and work it out.